OpenStudy (anonymous):

3x+2y=-19 x-12y=19 solve this for me anyone please :)

Directrix (directrix):

3x+2y=-19 x-12y=19 --> I am not following why there are two equality symbols in this. Would you check to see if you wrote it corrrectly? Thanks.

OpenStudy (anonymous):

no its right

OpenStudy (anonymous):

3x+2y=-19 x-12y=19

OpenStudy (mathstudent55):

3x + 2y = -19 x - 12y = 19 Solve the second equation for x: x = 12y + 19 Now substitute what x is equal to in first equation: 3(12y + 19) + 2y = -19 36y + 57 + 2y = -19 38y = -76 y = -2 Now substitute -2 in for y in first original equation: 3x + 2(-2) = -19 3x - 4 = -19 3x = -15 x = -5 Solution: x = -5, y = -2 This is the substitution method of solving a system of equations.

OpenStudy (anonymous):

thank you so much i appeciate it so much :)

OpenStudy (mathstudent55):

you're very welcome

OpenStudy (anonymous):

do you what 4x+y=8 -3x-y=0 is?

OpenStudy (anonymous):

or x+y=10 -x-2y=-14

OpenStudy (mathstudent55):

4x + y = 8 -3x - y = 0 This is a good system of equations to solve by the elimination method. With the elimination method, you add the two equations and eliminate one variable. Here notice that you have y in the first equation and -y in the second equation. By adding equations, y will be eliminated, and you can solve for x. 4x + y = 8 -3x - y = 0 ---------------- (add) x = 8 Now substitute the value of x, 8, in the first original equation and solve for y: 4(8) + y = 8 32 + y = 8 y = 8 - 32 y = -24 Solution: x = 8, y = -24

OpenStudy (mathstudent55):

The following system is also good for the elimination method. Add the two equations and eliminate x. Then solve for y. Once you have the value of y, plug it in the first equation and solve for x. x + y = 10 -x - 2y = -14